OpenStudy (anonymous):

how to simplify 4cos^2 x ???? :(

OpenStudy (anonymous):

uhm... can't be simplified any further

OpenStudy (anonymous):

yes :P \[4 \left(\begin{matrix}1+\cos2x \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

But I dont know how to arrive to this answer.. I dont know what to do!

OpenStudy (anonymous):

that's not simplification. It's just a different but equivalent form

OpenStudy (anonymous):

if you want to get rid of the power, i guess you could do \[\cos^2(x) = (1 + \cos^2(x))/2\]

OpenStudy (anonymous):

its the power reduction formula for cos^2(x)

OpenStudy (anonymous):

btw, that can be simplified to 2 + 2cos(2x)

OpenStudy (anonymous):

but why the 2 is not an exposant anymore ??

OpenStudy (anonymous):

that's wrong, its supposed to be\[\cos^2(x)=(1+\cos(2x))/2\]

OpenStudy (anonymous):

its the power reduction formula

OpenStudy (anonymous):

Thanks youuuuu a lot :)!!